OpenStudy (anonymous):

Convert to a fraction in lowest terms. 79.2%

OpenStudy (anonymous):

Since there are three digits, put it over 1000. 792/1000. Now simplify. 792/1000 = 396/500 = 198/250 = 99/125. That's as far as you can reduce it. You know it's the same value because if you divide it, you end up with 79.2%.

OpenStudy (anonymous):

thank you